# PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis

### v1

@@ 6 Conclusion
We propose a novel physics-aware dataset, PhysGaia, specifically designed to understand physics in dynamic scenes, particularly for Dynamic Novel View Synthesis (DyNVS). Comprising 17 diverse scenes, our dataset captures complex multi-body interactions with a wide variety of materials. Each scene is faithfully generated using material-specific physics solvers, ensuring adherence to physical laws and providing rich ground-truth physics data including 3D particle trajectories and physics parameters. This ground truth data uniquely enables the evaluation of physics reasoning. We also test state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and highlighting significant potential for improvement. We believe PhysGaia will be a critical resource that accelerates progress in physics-aware dynamic scene understanding.
Appendix A Details of PhysGaia
A.1 Scene Lists
Our dataset consists of 17 scenes divided into four categories: liquid, gas, viscoelastic substances, and textile. Each category contains 4 to 5 scenes, and the detailed scene names are listed in Table 5.
Table 5: List of scenes included in our PhysGaia dataset
Liquid Gas Viscoelastic substance Textile
@@ A.1 Scene Lists
A.2 Camera Trajectories
@@
Figure 7: Visualization of training camera’s traj.
Figure 7 visualizes our training camera trajectories, where we adopt either 2 or 4 cameras depending on the properties of each scene. For the cow, single flag, and tube scenes–which resemble $180^{\circ}$ settings with background walls placed behind the objects–we use only 2 cameras, as shown in Figure 7(a). In contrast, for $360^{\circ}$ scenes, we place 4 cameras at evenly spaced viewpoints around the scene, enabling full coverage from four directions. All cameras follow spiral trajectories to capture diverse views over time. For the monocular setup, we select one of the training cameras to generate the training data.
A.3 Implementation Details
All simulations were conducted on 4 NVIDIA Titan Xp GPUs and Intel Xeon E5-2630 v4 CPUs with 40 cores. While most scenes rendered in 2–3 hours, the lucy and hanok scenes required 12–24 hours due to their complex geometry and increased simulation costs. To supplement Section 4, we provide additional detailed simulation implementation along with the value of physics parameters. These parameters can be served as ground-truth labels for downstream physical reasoning tasks.
Liquid
We assume incompressible, non-Newtonian fluids in our simulations. Table 6 summarizes the density and viscosity values, which are the most important factors governing fluid motion. External forces are limited to gravity. For the ship scene, we reduced the particle separation from 0.1 to 0.05 to decrease initial particle displacement and improve simulation accuracy. Also, the particle separation during dynamic simulation is set to 0.1 for the cereal scene and 0.05 for the ship scene.
Table 6: Physics parameters used for liquid materials.
@@ A.3 Implementation Details
Gas
In gas simulations, temperature is the most critical physical parameter, governing buoyancy, expansion, and flow dynamics. We initialized the temperature to 3000K for gas materials. As temperature evolves locally due to gas movement and interactions, its spatiotemporal variation is represented as a temperature field defined on a voxel grid. These temperature fields are accessible via the provided simulation source file with Python API, similar to how flow fields are accessed for generating ground-truth trajectory data.
Viscoelastic substances
Gravity is applied to all scenes; however, for the cow scene, we introduce additional internal forces to induce cow motion and increase the complexity of the physical reasoning scenarios. Boundaries are set to be open in all directions, allowing objects to move freely without collisions against invisible walls. Table 7 lists detailed physics parameters, including Young’s modulus ($E$) and Poisson’s ratio ($P$), used in the simulations.
Table 7: Physics parameters used for viscoelastic materials.
Scene name Object Type $E$ $P$ Viscosity
@@
Cow Cow Elastic $5.0\times 10^{4}$ 0.45 -
Textile
For textiles, we primarily adopt the default configuration from the Vellum solver, with detailed modifications summarized in Table 8.
Table 8: Physics parameters used for textile materials.
@@
Flags Flag1 (small) $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
Flag2 (big) $1.0\times 10^{11}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
[…]
Cloth $1.0\times 10^{13}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
[…]
Lucy Cloth $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
Tube Tube flag $101.0\times 10^{10}$×$1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-2}$
A.4 Scene Visualization
Figure 8 and Figure 9 visualize more scenes in PhysGaia to supplement the Figure 1 in our main paper. As shown in Table 2 of the main paper, textile materials remain underexplored in physics-based datasets. To facilitate progressive research development in this domain, we include few simpler scenes such as single flag and tube that employ with only basic wind interactions. We believe this approach effectively supports advancements by providing more accessible starting points. Similarly, our dataset also includes few simpler gas scenes, simple smoke and falling, for the same reason as with textile scenarios: to facilitate progressive development. For the cow scene, internal forces intrinsic to the cow itself introduce unique and more complex dynamics, rather than using only gravity.
Figure 8: Examples from the proposed physics-aware dataset, PhysGaia.
(a) Gas: box-smoke (top-left), pisa (top-right)), simple smoke (bottom-left), and falling (bottom-right)
(b) Viscoelastic substances: pancake (top-left), jelly party (top-right), bouncing balls (bottom-left), and cow (bottom-right)
(c) Textile: lucy (top-left), basin (top-right), flags (mid-left), tube (mid-right), and single flag (bottom)
Figure 9: Examples from the proposed physics-aware dataset, PhysGaia. They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, viscoelastic substance, and textile. This dataset will foster physics reasoning in dynamic scenes.
Appendix B Details of Experimental Results
To supplement Table 3, Table 4, and Figure 5 of the main paper, we report more detailed quantitative results along with additional qualitative results.
B.1 Performance breakdown: monocular setting
We provide per-scene breakdown performance of monocular video reconstruction setting in Table 9. We report the performance of D-3DGS [36], 4DGS [37], STG [38], and SOM [15], which serve as the most common and recent baselines for the DyNVS task.
Table 9: Per-scene breakdown results for all 17 scenes under the monocular setting.

### v2

@@ 6 Conclusion
We propose a novel physics-aware dataset, specifically designed to understand physics in dynamic scenes, particularly for DyNVS. Our benchmark captures complex multi-body interactions with a wide variety of materials. Each scene is faithfully generated using material-specific physics solvers, ensuring adherence to physical laws and providing rich ground-truth physics data including 3D flows and configurations. This ground-truth data uniquely enables the evaluation of physical reasoning. We also test several tasks on PhysGaia, revealing their fundamental limitations in achieving physical realism and handling multi-objects, which highlights significant potential for improvement.
Acknowledgements
This work was partly supported by Samsung Research, Samsung Electronics Co., Ltd., the National Research Foundation of Korea (NRF) grant [RS-2022-NR070855, Trustworthy Artificial Intelligence], and the Institute of Information & communications Technology Planning & Evaluation (IITP) grants [RS-2025-25442338, AI star Fellowship Support Program (Seoul National University); RS-2022-II220959 (No.2022-0-00959), (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making; No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] funded by the Korea government (MSIT).
A Preliminary: 4D Gaussian Splatting
A.1 Gaussian primitive
3D Gaussian Splatting has recently achieved real-time rendering with state-of-the-art quality on static scenes. It adopts an explicit representation of the 3D scene through a collection of Gaussian ellipsoids $\Gamma=\left\{\gamma_{1},...,\gamma_{K}\right\}$. Each primitive $\gamma_{k}$ is defined by an unnormalized 3D Gaussian kernel $\mathcal{G}_{k}(\bm{x})$, parameterized by its mean $\bm{\mu}_{k}$ and covariance $\bm{\Sigma}_{k}$:
$\mathcal{G}_{k}(\bm{x};\bm{\mu}_{k},\bm{\Sigma}_{k})\coloneqq\exp\left(-\frac{1}{2}(\bm{x}-\bm{\mu}_{k})^{\top}\bm{\Sigma}_{k}^{-1}(\bm{x}-\bm{\mu}_{k})\right),$ (A)
where $\bm{\mu}_{k}\in\mathbb{R}^{3}$ denotes the primitive’s center position, $\bm{\Sigma}_{k}\in\mathbb{R}^{3\times 3}$ is an anisotropic covariance matrix, and $\bm{x}\in\mathbb{R}^{3}$ represents an arbitrary spatial coordinate. Since $\bm{\Sigma}_{k}$ must remain positive semi-definite during optimization—a constraint that is difficult to enforce directly—it is instead factorized into a rotation matrix $\bm{R}_{k}$ and a scaling matrix $\bm{S}_{k}$:
$\bm{\Sigma}_{k}\coloneqq\bm{R}_{k}\bm{S}_{k}\bm{S}_{k}^{\top}\bm{R}_{k}^{\top}.$ (B)
Beyond these Gaussian parameters $(\bm{\mu}_{k},\bm{R}_{k},\bm{S}_{k})$, each primitive also requires an opacity value $\alpha_{k}\in[0,1]$ and a feature vector $\bm{f}_{k}\in\mathbb{R}^{d}$, which typically encodes RGB colors or spherical harmonic (SH) coefficients. Thus, a Gaussian primitive is fully specified as $\gamma_{k}\coloneqq(\bm{\mu}_{k},\bm{R}_{k},\bm{S}_{k},\alpha_{k},\bm{f}_{k})$.
A.2 Differentiable rasterization
Before rendering the Gaussian primitives $\Gamma$ into the image space, each 3D Gaussian kernel, $\mathcal{G}_{k}(\bm{x};\bm{\mu}_{k},\bm{\Sigma}_{k})$, is first projected onto the 2D image space, resulting in a 2D Gaussian kernel, $\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})$. Here, $\pi:\mathbb{R}^{3}\to\mathbb{R}^{2}$ denotes the projection from world coordinates to image space. In the projected Gaussian representation, $\bm{r}\in\mathbb{R}^{2}$ indicates a pixel location in an image, and the 2D mean $\bm{\mu}_{k}^{\pi}\in\mathbb{R}^{2}$ and covariance $\bm{\Sigma}_{k}^{\pi}\in\mathbb{R}^{2\times 2}$ are given by
$\bm{\mu}_{k}^{\pi}\coloneqq\pi(\bm{\mu}_{k})\quad\quad\text{and}\quad\quad\bm{\Sigma}_{k}^{\pi}\coloneqq\bm{J}\bm{W}\bm{\Sigma}_{k}\bm{W}^{\top}\bm{J}^{\top},$ (C)
where $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation and $\bm{W}$ is the world-to-camera transformation matrix. When rendering the primitives in $\Gamma$ from a target camera view, they are sorted according to their depth relative to the camera center. The color of a pixel $\bm{r}$ is then determined by $\alpha$-blending:
$\mathbf{\hat{I}}(\bm{r})\coloneqq\sum_{k=1}^{K}\omega^{\pi}_{k}(\bm{r})c(\bm{f}_{k},\bm{r}),$ (D)
where $\omega^{\pi}_{k}(\bm{r})$ denotes the contribution of primitive $\gamma_{k}$ to pixel $\bm{r}$, and $c(\bm{f}_{k},\bm{r})$ is its corresponding color. If $\bm{f}_{k}$ encodes spherical harmonics coefficients, the color is evaluated using the view direction associated with $\bm{r}$; otherwise, $\bm{f}_{k}$ can directly represent the primitive’s RGB values. Following the $\alpha$-blending formulation in 3DGS [31], the visibility weight is defined as
$\omega^{\pi}_{k}(\bm{r})\coloneqq\alpha_{k}\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})\prod_{j=1}^{k-1}\left(1-\alpha_{j}\mathcal{G}^{\pi}_{j}(\bm{r};\bm{\mu}_{j}^{\pi},\bm{\Sigma}_{j}^{\pi})\right),$ (E)
where $\alpha_{k}\mathcal{G}^{\pi}_{k}(\bm{r};\bm{\mu}_{k}^{\pi},\bm{\Sigma}_{k}^{\pi})$ represents the opacity of the $k^{\text{th}}$ projected primitive at pixel $\bm{r}$, and the product term encodes the transmittance, i.e., how much light passes through the preceding primitives along the ray. Further details can be found in the original Gaussian Splatting paper [31].
A.3 Deformation for Dynamic Modelings
To represent 4D scenes with Gaussian Splatting, recent methods [79, 28, 74, 42, 70, 38] learn time-dependent deformations that map canonical 3D Gaussian primitives to their posed states over time. At time $t$, the position, rotation, and scale of the $k$-th primitive are updated as
$(\bm{\mu}_{k}^{t},\bm{R}_{k}^{t},\bm{S}_{k}^{t})=\big(\bm{\mu}_{k}+\phi_{\mu}(\cdot,t),\;\phi_{r}(\cdot,t)\,\bm{R}_{k},\;\bm{S}_{k}+\phi_{s}(\cdot,t)\big),$ (F)
where the deformation functions $\phi_{\mu}(\cdot,t)$, $\phi_{r}(\cdot,t)$, and $\phi_{s}(\cdot,t)$ denote learnable fields that may take as input various combinations of canonical primitive attributes, spatial coordinates, or time embeddings, depending on the parameterization. The rotation update $\phi_{r}(\cdot,t)$ is implemented as a residual rotation, typically parameterized via an exponential map to ensure $\phi_{r}(\cdot,t)\in SO(3)$. These functions are commonly instantiated with MLPs, learnable control points [70, 38, 28], HexPlane [74], or low-order polynomials [42]. The resulting time-varying primitive is then given by $\gamma_{k}(t)\coloneqq(\bm{\mu}^{t}_{k},\bm{R}^{t}_{k},\bm{S}^{t}_{k},\alpha_{k},\bm{f}_{k})$.
B Scene Composition
Our benchmark consists of 17 scenes, categorized into four material types: liquids, gases, rheological materials, and textiles. Each category contains 4 to 5 scenes, as listed in Table A. This section details the solver properties for each material and provides key simulation details regarding multi-body interactions. Please refer to our supplemental document for other simulation configurations.
Table A: List of scenes included in our PhysGaia benchmark. Our benchmark consists of 17 scenes, categorized into four material types: liquids, gases, rheological materials, and textiles.
Liquid Gas Rheological substance Textile
@@ B Scene Composition
Figure A and Figure B visualize more scenes in PhysGaia to supplement the Figure 3 in our main paper. As shown in Table 1 and Table 2 of the main paper, textile materials remain underexplored in physics-based datasets. To facilitate progressive research development in this domain, we include few simpler scenes such as single flag and tube that employ with only basic wind interactions. We believe this approach effectively supports advancements by providing more accessible starting points. Similarly, our benchmark also includes few simpler gas scenes, simple smoke and falling, for the same reason as with textile scenarios: to facilitate progressive development. For the cow scene, internal forces intrinsic to the cow itself introduce unique and more complex dynamics, rather than using only gravity.
(b) Gas: box-smoke (top-left), pisa (top-right)), simple smoke (bottom-left), and falling (bottom-right)
Figure A: Examples from the proposed physics-aware dataset, PhysGaia. They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, viscoelastic substance, and textile. This dataset will foster physics reasoning in dynamic scenes.
(a) Rheological substances: pancake (top-left), jelly party (top-right), bouncing balls (bottom-left), and cow (bottom-right)
(b) Textile: lucy (top-left), basin (top-right), flags (mid-left), tube (mid-right), and single flag (bottom)
Figure B: Examples from the proposed physics-aware benchmark, PhysGaia. They exhibit complex physical interactions between multiple bodies composed of diverse materials such as liquid, gas, rheological substance, and textile. This benchmark will foster physics reasoning in dynamic scenes.
C Common Setup
Simulator selection
Rendering
All frames are rendered at a resolution of $640\times 720$. We use path tracing with 256 samples per pixel and apply NVIDIA OptiX for denoising. Scenes are illuminated with 1–3 point lights (intensity: 600–4000); shadows are disabled in textile-focused scenes like tube-flag to emphasize geometry. All cameras follow spiral trajectories to capture diverse views over time. All simulations were conducted on 1 NVIDIA Titan Xp GPUs and Intel Xeon E5-2630 v4 CPUs with 40 cores. While most scenes rendered in 2–3 hours, the lucy and hanok scenes required 12–24 hours due to their complex geometry and increased simulation costs.
Camera Trajectories
@@ C Common Setup
Figure C: Visualization of training camera’s traj. All cameras follow spiral trajectories to capture diverse views over time. For the monocular setup, we select one of the training cameras to generate the training data.
Figure C visualizes our training camera trajectories, where we adopt either 2 or 4 cameras depending on the properties of each scene. For the cow, single flag, and tube scenes–which resemble $180^{\circ}$ settings with background walls placed behind the objects–we use only 2 cameras, as shown in Figure 3(a). In contrast, for $360^{\circ}$ scenes, we place 4 cameras at evenly spaced viewpoints around the scene, enabling full coverage from four directions. All cameras follow spiral trajectories to capture diverse views over time. For the monocular setup, we select one of the training cameras to generate the training data.
D Material-specific Physics Solver
D.1 Liquid
For liquid scenes, we adopt the Fluid-Implicit Particle (FLIP) solver [5] a hybrid particle-grid method. FLIP maintains particle velocities throughout the simulation, using the grid solely to compute and apply forces such as pressure and viscosity. This approach preserves fine-scale, high-frequency particle velocities, which is crucial for modeling realistic and rapid liquid behavior while adhering to the Navier–Stokes equations. While the Material Point Method (MPM) [66] can also model fluids, its direct velocity aggregation onto the grid limits its ability to capture highly dynamic fluid phenomena like splashing. Particle-only solvers [21] are another option, but FLIP is generally better suited for incompressible fluids due thanks to its hybrid grid-based representation.
Dynamic interactions
Physics law
We assume incompressible fluids in our simulations. This assumption simplifies the governing equations while maintaining high fidelity for typical liquid behaviors. Fluid motion is governed by the conservation of momentum and mass. We consider gravity as the sole external force. The momentum equation for our fluid, describing the conservation of momentum, is given by
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}\right)=-\nabla p+\mu\nabla^{2}\mathbf{u}+\rho\mathbf{g},$ (G)
where $\rho$ is density, $\mathbf{u}$ is velocity, $t$ is time, $p$ is pressure, $\mu$ is dynamic viscosity, and $\mathbf{g}$ is gravitational acceleration. The incompressibility condition ensures the conservation of mass, stating that the fluid’s volume remains constant as follows:
$\nabla\cdot\mathbf{u}=0,$ (H)
where $\mathbf{u}$ is velocity.
Physics parameters
Table B summarizes the density ($\rho$) and viscosity ($\mu$) values, which are the most important factors governing fluid motion. For the ship scene, we reduced the particle separation from 0.1 to 0.05 to decrease initial particle displacement and improve simulation accuracy. Also, the particle separation during dynamic simulation is set to 0.1 for the cereal scene and 0.05 for the ship scene.
Table B: Physics parameters used for liquid materials.
@@ D.1 Liquid
D.2 Gas
Dynamic interactions
Physics law
Our gas simulations, similar to liquids, are governed by the Navier-Stokes equations. However, we specifically account for the effects of buoyancy as an additional external force due to temperature differences.
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}\right)=-\nabla p+\mu\nabla^{2}\mathbf{u}+\rho(\mathbf{g}+\mathbf{f}_{b}),$ (I)
where $\mathbf{f}_{b}$ is the buoyancy force. The buoyancy force is derived from temperature differences relative to the ambient environment as follows:
$\mathbf{f}_{b}=k_{B}\cdot(T-T_{\text{ambient}})\cdot\mathbf{d}_{B},$ (J)
where $k_{B}$ is the buoyancy constant, $T$ is the local gas temperature, $T_{\text{ambient}}$ is the ambient temperature, and $\mathbf{d}_{B}$ is the buoyancy direction.
Physics parameters
In gas simulations, temperature is the most critical physical parameter, as it directly governs buoyancy, expansion, and overall flow dynamics. We initialized the temperature of gas materials to 3000K. As temperature evolves locally due to gas movement and interactions, its spatiotemporal variation is represented as a temperature field defined on a voxel grid. These temperature fields are accessible via the provided simulation source file with a Python API, similar to how flow fields are accessed for generating ground-truth trajectory data.
D.3 Rheological Substances
We use MPM [66] for rheological substances like snow and jelly. As an extension of FLIP [5], MPM is ideal for simulating chunk-based materials. It aggregates particle information onto a grid, performs computations, and then reprojects results back to particles, effectively capturing deformation and internal force propagation.
Dynamic interactions
Physics law
Simulating viscoelastic materials requires modeling both their elastic (solid-like) and viscous (fluid-like) responses. Our approach follows fundamental conservation laws while adopting simplified constitutive models commonly used in MPM-based visual simulations, namely linear elasticity and Kelvin–Voigt viscosity. The conservation of mass is described by the continuity equation as follows:
$\frac{D\rho}{Dt}+\rho(\nabla\cdot\mathbf{v})=0,$ (K)
where $\rho$ is the material density and $\mathbf{v}$ is the velocity field. Conservation of linear momentum is given by
$\rho\frac{D\mathbf{v}}{Dt}=\nabla\cdot\bm{\sigma}+\rho\mathbf{g},$ (L)
where $\bm{\sigma}$ is the Cauchy stress tensor and $\mathbf{g}$ is gravitational acceleration. We adopt a Kelvin–Voigt-type viscoelastic model, where the total stress is decomposed into elastic and viscous components as follows:
$\bm{\sigma}=\bm{\sigma}_{\text{elastic}}+\bm{\sigma}_{\text{viscous}}.$ (M)
The elastic stress is modeled using linear isotropic elasticity as follows:
$\bm{\sigma}_{\text{elastic}}=\frac{E\nu}{(1+\nu)(1-2\nu)}(\text{tr}(\bm{\epsilon}))\mathbf{I}+\frac{E}{2(1+\nu)}\bm{\epsilon},$ (N)
where $E$ is Young’s modulus, $\nu$ is Poisson’s ratio, and $\bm{\epsilon}$ is the strain tensor defined as $\bm{\epsilon}=\frac{1}{2}(\nabla\mathbf{u}+(\nabla\mathbf{u})^{T})$, where $\mathbf{u}$ is the displacement field. The viscous stress accounts for rate-dependent deformation as follows:
$\bm{\sigma}_{\text{viscous}}=2\mu\left(\mathbf{E}-\frac{1}{3}(\text{tr}(\mathbf{E}))\mathbf{I}\right)+\zeta(\text{tr}(\mathbf{E}))\mathbf{I},$ (O)
where $\mu$ is the dynamic viscosity and $\zeta$ is the bulk viscosity. The strain-rate tensor $\mathbf{E}$ is defined as $\mathbf{E}=\frac{1}{2}(\nabla\mathbf{v}+(\nabla\mathbf{v})^{T})$.
Physics parameters
Gravity is applied to all scenes; however, for the cow scene, we introduce additional internal forces to induce cow motion and increase the complexity of the physical reasoning scenarios. Boundaries are set to be open in all directions, allowing objects to move freely without collisions against invisible walls. Table C lists detailed physics parameters, including Young’s modulus ($E$) and Poisson’s ratio ($\nu$), used in the simulations.
Table C: Physics parameters used for rheological materials.
Scene name Object Type $E$ $\nu$ Viscosity
@@ D.3 Rheological Substances
Bouncing balls Balls Elastic $1.0\times 10^{5}$ 0.50 -
Fishbowl Static - - -
Jelly party Jelly Elastic $8.0\times 10^{4}$ 0.45 -
Cow Cow Elastic $5.0\times 10^{4}$ 0.45 -
D.4 Textile
Dynamic interactions
To simulate interactions between objects and textiles–where both move and exchange forces, as seen particularly in the basin scene–we employ the shape match constraint on objects. This constraint helps maintain the overall shape of rigid objects by driving points toward their rest configuration, allowing the material to preserve its structural integrity while still interacting dynamically with textiles and other objects. For the lucy scene, we increased the simulation sub-step count fivefold over the default to robustly handle collisions with the statue’s complex geometry. To simulate wind-induced fluttering in the flags, single-flag, and tube scenes, we applied external forces using a POP Wind node that blows parallel to the ground plane.
Physics law
We simulate textile materials using the Vellum solver [63], which is based on Extended Position-Based Dynamics (XPBD) [50]. XPBD extends classical Position-Based Dynamics (PBD) by introducing a compliance parameter, allowing constraint stiffness to be decoupled from the time step and iteration count. Unlike force-based methods, XPBD does not explicitly solve equations of motion. Instead, particle positions are updated by iteratively projecting them to satisfy geometric constraints. The core algorithm for each simulation substep is shown in Algorithm A.
Algorithm A Extended Position-Based Dynamics (XPBD) Algorithm (per substep)
1:Input:
@@ D.4 Textile
[…]
6: Iterations: Number of constraint projection iterations
7:// 1. Predict positions
8:for each particle $i$ do
9:  $\mathbf{p}^{*}_{i}\leftarrow\mathbf{p}_{i}+\Delta t\mathbf{v}_{i}+\Delta t^{2}\mathbf{a}_{\text{ext},i}$
10:end for
11:// 2. Constraint projection
12:for $k=1$ to Iterations do
13:  for each constraint $C_{j}$ do
14:   compute $\Delta\mathbf{p}_{i}$ using Eq. (P)
15:   $\mathbf{p}^{*}_{i}\leftarrow\mathbf{p}^{*}_{i}+\Delta\mathbf{p}_{i}$
16:  end for
17:end for
18:// 3. Velocity update
19:for each particle $i$ do
20:  $\mathbf{v}_{i}\leftarrow(\mathbf{p}^{*}_{i}-\mathbf{p}_{i})/\Delta t$
21:  $\mathbf{p}_{i}\leftarrow\mathbf{p}^{*}_{i}$
22:end for
We primarily employ stretch and bend constraints to model textile behavior. For a constraint $C(\mathbf{p})$, where $C(\mathbf{p})$ is a scalar constraint function that measures the violation of a geometric constraint, such as maintaining distances (stretch) or bending angles between particles, with $C(\mathbf{p})=0$ indicating that the constraint is satisfied, the position correction is as follows:
$\Delta\mathbf{p}_{i}=-\Delta\lambda\,\mathbf{w}_{i}\nabla_{\mathbf{p}_{i}}C,$ (P)
where
$\Delta\lambda=\frac{C(\mathbf{p})}{\sum_{j}\mathbf{w}_{j}|\nabla_{\mathbf{p}_{j}}C|^{2}+\tilde{\alpha}},\quad\tilde{\alpha}=\frac{\alpha}{\Delta t^{2}}.$ (Q)
Here, $\alpha$ denotes the compliance parameter (inverse stiffness).
Physics parameters
For textiles, we primarily adopt the default configuration from the Vellum solver, with detailed modifications summarized in Table D. In addition to stiffness, we specify a damping ratio for each constraint to reduce oscillations and stabilize the simulated cloth.
Table D: Physics parameters used for textile materials.
@@
Flags Flag1 (small) $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
Flag2 (big) $\times 10^{11} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
Basin Boxes $\times 10^{10} 1.0×10$$\times 10^{-3}$ N/A N/A
Cloth $\times 10^{13} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
Single flag Flag $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
Lucy Cloth $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
Tube Tube flag $\times 10^{10} 1.0×10$$\times 10^{-3}$ $\times 10^{-4} 1.0×10$$\times 10^{-2}$
E Details of Experimental Results
E.1 Implementation details
We implement recent DyNVS methods, including D-3DGS [79], 4DGS [74], STG [42], MoSca [38], and SOM [70]. We follow the default training settings provided for each method. For 4DGS, we adopt the training configuration used for the HyperNeRF [56] dataset and reduce the grid learning rate to improve training stability. For SOM, we apply the recommended hyperparameters for the DyCheck [18]dataset. For point cloud initialization, we run COLMAP with dense matching and fusion, followed by uniform downsampling to approximately 40,000 points. All experiments are conducted on NVIDIA RTX A5000 and A6000 GPUs, with training times ranging from 30 minutes to 2 hours depending on the method. Note that all experiments are implemented based on their public codes  [^4] D-3DGS: https://github.com/ingra14m/Deformable-3D-Gaussians [^5] STG: https://github.com/oppo-us-research/SpacetimeGaussians6664DGS: https://github.com/hustvl/4DGaussians [^7] SOM: https://github.com/vye16/shape-of-motion [^8] MoSca: https://github.com/JiahuiLei/MoSca.
E.2 Average performance: all scenes
Table E presents the average performance of all methods under both monocular and multiview settings, aggregated across all 17 scenes. For SOM [70], we observed better performance when scaling the estimated depth map using the COLMAP point cloud obtained from dense matching, compared to sparse matching.
Table E: Average quantitative results for both monocular and multiview settings, averaged across all 17 scenes. While multiview setups generally offer better reconstruction performance than monocular ones, even multiview results achieve PSNR scores below 30. This highlights the substantial difficulty in reconstructing the complex multi-body interactions in our benchmark.
Method Monocular Multiview
D-3DGS [79] 21.7 0.86 0.18 24.2 0.89 0.14
4DGS [74] 22.7 0.85 0.19 24.4 0.87 0.17
STG [42] 19.3 0.76 0.30 21.0 0.79 0.30
MoSca [38] 19.5 0.76 0.34 N/A N/A N/A
SoM [70] 19.3 0.80 0.26 N/A N/A N/A
E.3 Performance breakdown: monocular setting
We provide per-scene breakdown performance of monocular video reconstruction setting in Table F. We report the performance of D-3DGS [79], 4DGS [74], STG [42], MoSca [38] and SOM [70], which serve as the most common and recent baselines for the DyNVS task.
Table F: Per-scene breakdown results for all 17 scenes under the monocular setting.